At present, Lester's age is 30% of his mother, Lisa's age. Twenty years from now, his age will be 58% that of his mother's age. How old is Lester?
Answer by macston(5194)

L=Lester's age ; M=mother's age
L=0.30M; L+20=0.58(M+20) Substitute for L in second equation
0.30M+20=0.58(M+20)
0.30M+20=0.58M+11.6 Subtract 0.30M from each side
20=0.28M+11.6 Subtract 11.6 from each side.
8.4=0.28M divide each side by 0.28
30=M Mother's age is 30
Lester is 30% mother=(0.30)(30 years)=9 years old ANSWER Lester is 9 years old
CHECK
In 20 yrs Lester will be 58% his mother's age
L+20=0.58(M+20)
29=0.58(30+20)
29=0.58(50)
29=29